# Asymmetric Effects Underlying Dynamic Heterogeneity in Miscible Blends of Poly(methyl methacrylate) with Poly(ethylene oxide)

### v1

@@ 3.3 Influence of local environment on segmental mobility
Figure 2 shows that the influence of local environment is asymmetric between species and strongly temperature-dependent. In cases where $\mu_{i,\Delta t}$ is enhanced upon blending (above the dashed line), $\mu_{i,\Delta t}$ gradually approaches the neat reference, $\mu_{i,\Delta t}^{\circ}$, as the local environment becomes enriched in the same species; the notion of gradual in this context reflects a vanishing of the first derivative, $\frac{\partial\mu_{i,\Delta t}}{\partial\tilde{\phi}_{i}}$. Where $\mu_{i,\Delta t}$ is suppressed (below the dashed line), $\mu_{i,\Delta t}$ approaches $\mu_{i,\Delta t}^{\circ}$ more abruptly. While all dynamics tend to the neat reference in the limit of that the local environment is enriched in that species, these behaviors, which are species-agnostic, reveal that suppression effects are more readily evident than enhancements. In other words, the local coupling of polymers with the opposing species has a much stronger magnitude of effect of reducing the mobility of faster-moving chains than enhancing the mobility of slower-moving chains in a blend. Upon investigation of the behavior of $\mu_{i,\Delta t}-\mu_{i,\Delta t}^{\circ}$ normalized by $\mu_{i,\Delta t}^{\circ}$ shown in Figure S3, it is evident, however, that the relative change in $\mu_{i,\Delta t}$ is actually much larger for species with enhanced mobility than suppressed mobility.
@@ 3.4 Rouse mode analysis of collective dynamics in the melt state
To gain insight into how blending affects the length scale of altered collective dynamics of PEO and PMMA, we perform a Rouse mode analysis on chains of both species. The departure from local dynamics to collective dynamics allows us to probe longer time and length scale behaviors in the blends. This analysis is limited to systems where all species remain in the melt state, both given the computational challenges of equilibration and statistical convergence as well as the dramatic increase in relaxation times expected at temperatures below the $T_{\text{g}}$ of PEO or PMMA. Results are shown at the highest temperature of 500 K in Figure 3 and at 360 K in the Supplementary Information, Figure S6.
Figure 3: Rouse mode analysis at 500 K for chains in blends ov varying composition. The effective Rouse relaxation time $\tau^{\text{eff}}_{p}$ as a function of sub-chain length $N/p$ for (A) PEO and (B) PMMA. The mode-dependent stretching exponent $\beta_{p}$ for (C) PEO and (D) PMMA. For $p>8$, symbols for data are only shown for every third value of $p$ for visual clarity. Dashed black lines in panels (A) and (B) are a guide to the eye to indicate the expected ideal scaling of $\tau_{p}\sim p^{-2}$. The position of the line is the same across panels and is set to align with the behavior of neat PEO. Results for chains in blends with the most extreme compositions are outlined in black for visual clarity of trends. Error bars reflect standard errors from three independent systems and are generally smaller than the symbol size.
@@
Figure 3A shows that the characteristic relaxation time $\tau^{\text{eff}}_{p}$ in PEO/PMMA blends deviates from that of pure PEO (dashed black line) across all blend compositions and for all mode numbers beyond the monomer scale ($N/p>1$). Pure PEO approximately follows the expected Rouse scaling, $\tau_{p}\sim p^{-2}$,74 when using $\tau^{\text{eff}}_{p}$, which reflects stretched rather than single-exponential relaxation. As the mode number increases ($p\to N$), $\tau^{\text{eff}}_{p}$ for all blends converge, suggesting that relaxation behavior at the monomer scale, for PEO, is largely unaffected by the surrounding melt environment. In contrast, deviations from pure PEO behavior become more pronounced at lower $p$ as the blend becomes more PMMA-rich (decreasing $x^{\text{(PEO)}}$), indicating altered relaxation at longer length scales.
Figure 3B shows that the relaxation dynamics of PMMA chain segments, across all values of $N/p$, increasingly resemble those of PEO as $x^{\text{(PEO)}}$ increases. A notable distinction between PMMA and PEO is the non-monotonic behavior of $\tau^{\text{eff}}_{p}$ for PMMA. Relaxation times begin to increase around $N/p=2$, likely reflecting the dominance of local PMMA environments at short sub-chain lengths. This results in a nearly composition-independent enhancement of dynamics at short length scales, followed by growing deviations from neat PMMA behavior at longer scales. In contrast to PEO, where sufficiently short sub-chains recover the relaxation behavior of neat PEO, short PMMA sub-chains in blends never fully recover neat PMMA behavior. This highlights the persistent and composition-wide influence of PEO on PMMA dynamics, with the effect becoming more pronounced at larger length scales. The enhancement of PMMA dynamics due to PEO is also much stronger than the suppression of PEO dynamics due to PMMA.
Figure 3C shows that the enhanced deviation of PEO relaxation from the neat-polymer dynamics is accompanied by stronger departures from ideal Rouse-like behavior. Specifically, the stretching exponent $\beta_{p}$ quantifies the extent to which the $p$th mode resembles ideal Rouse dynamics, for which $\beta_{p}=1$. As $x^{\text{(PEO)}}$ decreases, $\beta_{p}$ decreases across all modes, indicating increasingly non-ideal relaxation behavior across all segment length scales. Additionally, for a given blend composition, relaxation behavior becomes increasingly non-ideal at shorter length scales (lower $N/p$). Conversely, $\beta_{p}$ appears to plateau for $N/p\geq 7$ and decrease for $N/p<7$, except for neat PEO, which only exhibits a decrease in $\beta_{p}$ for $N/p<2$. In addition, in the $N/p<7$ regime, the degree of deviation strongly depends on blend composition. Together, these results suggest that dynamical coupling of the blend environment emerges at length scales of around 7 monomers of PEO and reflect the sensitivity of collective relaxation to compositional heterogeneity.
Analysis of the stretching coefficients for PMMA in Figure 3D further highlight PEO as the dynamically dominant component in the blend on longer time scales ($\geq 2000$ ps). While neat PMMA is strongly non-ideal in terms of its relaxation dynamics, the dynamics of PMMA in blends become more Rouse-like with increasing PEO content. Like PEO, PMMA exhibits enhanced deviation from ideality at shorter segment lengths; however, the onset of this deviation is composition-dependent for PMMA. In PEO-rich blends, the transition to enhanced non-ideality for PMMA begins around sub-chains of five monomers, which is comparable to the seven-monomer threshold observed for PEO in blends (Figure 3C). Conversely in PMMA-rich systems, $\beta_{p}$ decreases for sub-chains of about two monomers. Although these conclusions are derived from results at 500 K, they extend to systems in the melt state at 360 K, as shown in the Supplementary Information, Figure S6.
@@ 3.5 Interplay Between Local Mobility and Relaxation Timescales
At this stage, we remark on an apparent contradiction between the results shown in Figures 3A and B and those in Figures 2C and 2F. Figure 2 indicates that PEO segmental mobility is suppressed even in the presence of a small number of nearby PMMA segments, whereas PMMA mobility is only notably affected when the local environment becomes significantly enriched in PEO ($\tilde{\phi}_{i}^{\text{(PMMA)}}$ < 0.5). In contrast, the Rouse mode analysis suggests a stronger overall influence of PEO on PMMA than the reverse. In addition, the relaxation times for PMMA do not converge to the neat reference value, even for short sub-chains (high $p$), whereas PEO dynamics do recover in similar limits.
Two factors help reconcile this discrepancy. First, although PEO appears more affected in absolute terms in Figure 2, normalizing by the neat-polymer reference reveals that PMMA experiences a much larger relative enhancement in dynamics in PEO-rich environments, as shown in the Supplementary Information, Figure S3. Second, there is also a fundamental distinction between the two measured quantities. Figure 2 relates to a segmental mobility, based on short-time fluctuations over a 100 ps interval, and Figure 3 relates to segmental relaxation, which reflects structural decorrelation, which occurs at much longer timescales. Thus, while local dynamics may appear similar over short times, this does not necessarily imply similar behavior in longer-time relaxation. We suspect that results between these analyses would begin to align if the time interval for assessing segmental mobility was substantially increased. This underscores a key nuance of dynamic heterogeneity, in that its effects depend on the timescale of the process being observed.
@@ 4 Conclusions
In this study, we systematically investigated blends of PEO and PMMA to uncover the molecular origins and consequences of dynamic heterogeneity at the nanoscale. Our study considered several facets, including the impact of overall blend composition, the influence of environmentally local composition, and the impact of temperature, assessing conditions above, between, and below the $T_{\text{g}}$ of pure PEO and PMMA. We further quantified dynamic heterogeneity based on effects of local segmental mobility or fluctuations as well as larger-scale collective relaxation behaviors via Rouse mode analysis. The collection of simulations and analyses together provides a more comprehensive view relative to prior contributions that may have focused on a more limited set of compositions or temperatures.
Differences in local structure were deduced to asymmetrically impact local segmental mobilities. In particular, when accounting for differences in $T_{g}$, PEO segmental dynamics were enhanced relative to neat PEO, whereas the segmental dynamics of PMMA in blends were overall similar to $T_{g}$-equivalent neat PMMA systems. This disparity was elucidated by free-volume effects, where PEO segmental fluctuations correlated well with increases in local free volume, while PMMA dynamics were less impacted by reductions in free volume. Changes in the local environment also led to asymmetric effects on local segmental mobilities at short times. In terms of absolute magnitudes, suppression of fluctuations due to a “slow” local environment was a stronger effect than enhancements by a “fast” local environment. However, relative to a neat-melt reference, the presence of PEO enhanced PMMA dynamics far more than PMMA suppressed those of PEO.
Rouse mode analysis for systems in the melt state further clarified the effects of compositional heterogeneity. Notably, PEO relaxation behavior approached that of the neat polymer in environments dominated by PEO (e.g., either sufficiently short length scales or in PEO-rich blends). In contrast, PMMA dynamics were consistently influenced by the presence of PEO, with PMMA segments in PEO-rich blends exhibiting PEO-like relaxation behavior except at the shortest length scales. Finally, the enhancement of PMMA relaxation dynamics in PEO-rich blends was substantially greater than the suppression of PEO dynamics in PMMA-rich blends, at least in the melt state. Together, these findings emphasize a directional coupling, wherein PMMA dominates the behavior of the blend at short timescales (100 ps) while PEO dynamics dominate at longer timescales ($\geq 2000$ ps).
These findings extend our molecular-level understanding of how blending dynamically dissimilar polymers alters both local and collective segmental motions. The observation that PEO dynamics recover under certain conditions, while PMMA dynamics remain broadly perturbed, highlights a fundamental asymmetry in dynamical influence. Whether these behaviors generalize to other polymer blends with disparate relaxation properties or other thermal conditions remains an open question. Future work may examine the transferability of these packing-structure-dependent effects to other systems and examine their implications for macroscopic properties.10, 75 In addition, our results suggest that one species may have a stronger influence on dynamics at short times, while another may dominate at longer times, and the influence of a species may be apparent only at certain degrees of local composition. This perspective may offer a strategy to tune certain properties (e.g., viscoelastic properties or diffusion rates of other components, like ions or penetrants) across multiple length scales by adjusting the content and morphology of dynamically distinct components.

### v2

@@ 3.3.1 Analysis of local mobility-composition coupling
Figure 2 shows that the influence of local environment is asymmetric between species and strongly temperature-dependent. The temperature-dependence is first apparent by comparing the results for PEO where $\mu_{i,\Delta t}$ is enhanced by blending at a temperature below both component $T_{\text{g}}$ (Figure 2A) whereas it is suppressed upon blending at higher temperatures (Figures 2B,C). Meanwhile, that the influence is asymmetric between species is evident by contrast with the PMMA results (Figures 2D-F), which display nominally the opposite behavior of enhancement upon blending at high temperatures and minor suppression upon blending at low temperatures. The observations at 360 K and 500 K are physically intuitive, as PEO-rich environments tend to exhibit larger mobilities, while PMMA-rich environments are slower. The observations at 220 K where mobilities appear enhanced in PMMA-rich environments are likely due to the free-volume effects described in Section 3.2.
The dependence of segmental mobility on local composition also displays intriguing differences between suppression and enhancement effects in blends. In cases where $\mu_{i,\Delta t}$ is enhanced upon blending (above the dashed line), $\mu_{i,\Delta t}$ gradually approaches the neat reference, $\mu_{i,\Delta t}^{\circ}$, as the local environment becomes enriched in the same species; the notion of gradual in this context reflects a vanishing of the first derivative, $\frac{\partial\mu_{i,\Delta t}}{\partial\tilde{\phi}_{i}}$. Where $\mu_{i,\Delta t}$ is suppressed (below the dashed line), $\mu_{i,\Delta t}$ approaches $\mu_{i,\Delta t}^{\circ}$ more abruptly. While all dynamics tend to the neat reference in the limit of that the local environment is enriched in that species, these behaviors, which are species-agnostic, reveal that suppression effects are more readily evident than enhancements. In other words, the local coupling of polymers with the opposing species has a much stronger magnitude of effect whereby the mobility of faster-moving chains is reduced more than that of slower-moving chains is enhanced in a blend. Upon investigation of the behavior of $\mu_{i,\Delta t}-\mu_{i,\Delta t}^{\circ}$ normalized by $\mu_{i,\Delta t}^{\circ}$ shown in Figure S5, it is evident, however, that the relative change in $\mu_{i,\Delta t}$ is actually much larger for species with enhanced mobility than suppressed mobility.
@@
3.3.2 Analysis of local composition distributions
While Figure 2 demonstrates how local composition leads to deviations in segmental dynamics from neat-polymer behavior, further understanding blend properties requires knowledge of how frequently such environments occur and how they are distributed. To this end, we characterize the local environments sampled by PEO and PMMA with the objective to relate these to deviations from $T_{\text{g}}$-equivalent neat dynamics. We compute probability distributions of local intermolecular PMMA volume fractions surrounding particles from PEO, $\phi_{\text{inter,PMMA}}^{\text{(PEO)}}$, and PMMA, $\phi_{\text{inter,PMMA}}^{\text{(PMMA)}}$, and characterize their breadth using a normalized Shannon entropy, $\tilde{H}$, where lower $\tilde{H}$ indicates more homogeneous environments and higher values indicates environments are more heterogeneously distributed.
Figure 3A shows how the distributions of intermolecular PMMA volume fractions around PEO units vary with blend composition at 500 K; analogous results at 220 K and 360 K are provided in the Supplementary Information (Figure S8). For PEO-rich blends, the distributions peak near $\phi_{\text{inter,PMMA}}^{(\text{PEO})}=0$, indicating that most PEO atoms experience negligible local PMMA content, consistent with the overall scarcity of PMMA. As $x^{(\text{PEO})}$ decreases and there is more PMMA present in the blend, the distributions shift toward higher $\phi_{\text{inter,PMMA}}^{(\text{PEO})}$. The distributions also broaden, revealing that local environments become increasingly heterogeneous as the blend composition becomes more balanced. This trend is reflected quantitatively in the normalized Shannon entropy, $\tilde{H}$ (Figure 3B), which rises with decreasing $x^{(\text{PEO})}$, signifying a wider range of local PMMA fractions sampled by PEO atoms. The entropy plateaus for $x^{(\text{PEO})}\lesssim 0.6$, indicating that the extent of heterogeneity becomes composition-independent once sufficient PMMA is present in the blend.
Figure 3: Characterization of local composition distributions around PEO. (A) Probability distributions of intermolecular PMMA volume fraction around PEO and (B) corresponding normalized Shannon entropies. Results are shown for systems at 500 K. The color gradient corresponds to a gradient in blend composition containing the most PEO (dark) to the least PEO (light). The dashed red line is provided as a guide to the eye.
Figure 4A shows how the distributions of intermolecular PMMA volume fractions around PMMA units vary with blend composition. As in Figure 3A, the distributions shift to higher $\phi_{\text{inter,PMMA}}^{(\text{PMMA})}$ with decreasing $x^{(\text{PEO})}$, reflecting the overall increase in PMMA content. However, unlike the distributions around PEO, their shapes remain comparatively narrow and composition-invariant, indicating that the local environments of PMMA change in a more uniform and predictable manner. This behavior is corroborated by the normalized Shannon entropy in Figure 4B, which remains nearly constant across compositions and is systematically lower than for PEO, signifying more homogeneous local surroundings. Together, these results highlight another asymmetry between PEO and PMMA upon blending. Namely, while the local environment of PMMA changes in a consistent way that can be tracked by the mean composition, as in the Lodge–McLeish framework,48 the local environment of PEO cannot because the nature of the distribution itself changes significantly with composition. In addition, PEO segments experience overall more local compositional heterogeneity.
Figure 4: Characterization of local composition distributions for PMMA. (A) Probability distributions of intermolecular PMMA volume fraction around PMMA and (B) corresponding normalized Shannon entropies. Results are shown for systems at 500 K. The color gradient corresponds to a gradient in blend composition containing the most PEO (dark) to the least PEO (light). The black dashed line in (A) is the distribution of neat PMMA. The red dashed line in (B) is provided as a guide to the eye.
We attribute the dynamical response of PEO in the blends to the increased heterogeneity and composition dependence of its local environment with respect to surrounding PMMA. These features correlate strongly with the deviations of local segmental mobilities from those of $T_{\text{g}}$-equivalent neat systems (Figure LABEL:figMSFrelt). For PEO, the growing diversity of local intermolecular PMMA environments with decreasing $x^{(\text{PEO})}$ (Figure 3A) coincides with its marked departure from the neat reference behavior. In contrast, the local environments of PMMA evolve more uniformly with composition (Figure 4), resulting in dynamics that follow expectations based on the overall reduction in $T_{\text{g}}$ with PEO incorporation. Distributions of local intramolecular self-contributions (Figure S6) are comparatively insensitive to composition, as expected. Meanwhile, the distributions of local intermolecular PEO contributions (Figure S7) again change simply in response to composition for PMMA, and for PEO, there is a modest increase in heterogeneity at low PMMA blending fractions, but this is less pronounced than the strong compositional diversity observed for surrounding PMMA (Figure S8). However, the distributions of local environments show minimal temperature dependence (Figures S6-S9), indicating that increased compositional heterogeneity alone does not account for the deviations observed well below $T_{\text{g}}$ (220 K). Instead, at such temperatures, the relative dynamics are better explained by differences in free-volume between the probed systems and their references (Figures LABEL:figMSFreltC,D).
@@ 3.4 Rouse mode analysis of collective dynamics in the melt state
To gain insight into how blending affects collective dynamics of PEO and PMMA, we perform a Rouse mode analysis of chains of both species. This approach shifts the focus from local segmental fluctuations to collective sub-chain dynamics, enabling us to probe relaxation across longer times and length scales. In this way, our aim is not to test whether the dynamics are strictly Rouse-like but rather use Rouse coordinates as a convenient collective-variable framework to assess how relaxation times vary with sub-chain length and composition. Given the computational challenges of equilibration and convergence as well as the dramatic increase in relaxation times expected at temperatures below $T_{\text{g}}$, our analysis is restricted to systems at 500 K (Figure 5).
@@
Figure 5A shows that the characteristic relaxation times, $\tau^{\text{eff}}_{p}$, of PEO in PEO/PMMA blends deviates from those of pure PEO, across all blend compositions and for all mode numbers. Pure PEO approximately follows the expected Rouse scaling (dashed black line), $\tau_{p}\sim p^{-2}$,88 albeit when using $\tau^{\text{eff}}_{p}$. As the mode number increases ($p\to N$), $\tau^{\text{eff}}_{p}$ for all blends begins to converge; this implies that the relaxation behavior of the smallest sub-chains of PEO are similar, whether in blends or in neat melts.
It should be noted that the classic Rouse scaling of $\tau_{p}\sim p^{-2}$ is only expected down to the Rouse bead size, which is distinct from the Kuhn length and can span multiple chemical monomers.89 Here, the apparent convergence in relaxation times occurs for sub-chains likely smaller than a canonical Rouse bead. Meanwhile, deviations from pure PEO behavior become more pronounced at lower $p$ as the blend becomes more PMMA-rich (decreasing $x^{\text{(PEO)}}$). Furthermore, since $\tau_{p}^{\mathrm{eff}}$ represents the mean relaxation time extracted from stretched exponential decays, apparent deviations from ideal scaling may reflect both physical departures90, 91, 92 from Rouse dynamics as well as non-exponential relaxation. All together, these results intuitively suggest that chain relaxation at shorter length scales is less affected by the surrounding melt environment than at longer length scales.
Figure 5B shows that variations in relaxation behavior of PMMA chain segments with $N/p$ are functionally similar across blend compositions, resulting in systematic vertical offsets in relaxation times. This indicates that blending primarily shifts the overall timescale of PMMA dynamics while preserving their scaling, suggesting that the data might be collapsed by a composition-dependent rescaling factor, governed by the distance from $T_{\text{g}}$ (see also Figure S15). Furthermore, the relaxation times for PMMA deviate from classical Rouse scaling across all mode numbers and compositions, exhibiting an upturn at short sub-chain lengths. This anomalous slowdown likely reflects that the relatively stiff PMMA chains contain only a few effective Rouse beads, such that the shortest modes are dominated by local constraints rather than entropic relaxation. This contrasts with Figure 5A, where PEO relaxation times scale differently with $N/p$ as a function of blend composition yet converge for short sub-chains, indicating that PEO dynamics can recover neat-like behavior at sufficiently small length scales. In comparison, PMMA dynamics remain uniformly perturbed by PEO across all scales.
Figure 5: Rouse mode analysis at 500 K for chains in blends of varying composition. The effective Rouse relaxation time $\tau^{\text{eff}}_{p}$ as a function of sub-chain length $N/p$ for (A) PEO and (B) PMMA. For $p>8$, symbols for data are only shown for every third value of $p$ for visual clarity. Dashed black lines are a guide to the eye to indicate the expected ideal scaling of $\tau_{p}\sim p^{-2}$. The position of the line is the same across panels and is set to align with the behavior of neat PEO. Results for chains in blends with the most extreme compositions are outlined in black for visual clarity of trends. Error bars reflect standard errors from three independent systems and are generally smaller than the symbol size. Transparent markers are used for $p$ corresponding to sub-chains equal to or less than an estimated Kuhn length. Deviations from the ideal $p^{-2}$ scaling are expected for modes shorter than the Rouse bead size, which is generally larger than a Kuhn length.
@@ 3.5 Interplay Between Local Mobility and Relaxation Timescales
At this stage, we remark on an apparent contradiction between the results shown in Figures 5A and 5B and those in Figures 2C and 2F. Figure 2 indicates that PEO segmental mobility is suppressed even in the presence of a small number of nearby PMMA segments, whereas PMMA mobility is only notably affected when the local environment becomes significantly enriched in PEO ($\tilde{\phi}_{i}^{\text{(PMMA)}}$ < 0.5). In contrast, the Rouse mode analysis suggests a strong overall influence of PEO on PMMA because relaxation times for PMMA do not converge to the neat reference, even for short sub-chains, whereas those for PEO do.
Two factors help reconcile this discrepancy. First, although PEO appears more affected in absolute terms in Figure 2, normalizing by the neat-polymer reference reveals that PMMA experiences a much larger relative enhancement in dynamics in PEO-rich environments (Figure S5). Second, there is also a fundamental distinction between the two measured quantities. Figure 2 relates to a segmental mobility, based on short-time fluctuations over a 100 ps interval, and Figure 5 relates to segmental relaxation, which reflects structural decorrelation, which occurs at much longer timescales. Thus, while local dynamics may appear similar over short times, this does not necessarily imply similar behavior in longer-time relaxation. We suspect that results between these analyses would begin to align if the time interval for assessing segmental mobility was substantially increased. This underscores a key nuance of dynamic heterogeneity, in that its effects depend on the timescale of the process being observed.
@@ 4 Conclusions
The seemingly straightforward dependence of PMMA dynamics on a shifted $T_{\text{g}}$ in contrast to the composition-dependent dynamics of PEO in PEO/PMMA blends have been the subject of several prior works.41, 42, 43, 40 Here, using atomistic molecular dynamics simulations, we extend these observations by systematically examining how overall and local composition influence segmental mobility and collective relaxation, across blend compositions, multiple length scales, and thermal regimes. Analyses of both local fluctuations and sub-chain relaxation dynamics reveal a coherent physical picture linking free-volume effects and local compositional heterogeneity to deviations from neat-polymer behavior.
Local environments were found to asymmetrically affect segmental mobilities. At short times, suppression by “slow” environments was stronger than enhancement by “fast” ones; nonetheless, PEO accelerated PMMA dynamics far more than PMMA suppressed PEO. Accounting for $T_{\text{g}}$ differences, PEO segments in blends exhibited enhanced dynamics relative to neat PEO, whereas PMMA segments in blends remained comparable to $T_{\text{g}}$-equivalent neat PMMA systems. These effects were directly connected to variations in local composition. PEO dynamics increased with local free volume, which is heterogenously distributed when blending with PMMA, yielding a response that differs from mean-field expectations. Meanwhile, PMMA dynamics varied more uniformly with blend composition, as does the local compositional variations of PMMA segments.
Rouse-mode analysis at 500 K revealed complementary signatures at collective length scales. PEO relaxation approached neat-like behavior at short sub-chain lengths or in PEO-rich blends, whereas PMMA relaxation was uniformly accelerated across all modes with increasing PEO content. This behavior suggests a more idealized, composition-dependent response for PMMA compared to the strongly heterogeneous PEO dynamics. In tandem with our observations on the local enhancement of PMMA dynamics in the vicinity of PEO, we suggest this arises from a nanoscale facilitation mechanism whereby locally softened PMMA near PEO propagates enhanced mobility to neighboring regions. This interpretation resonates with the dynamic coupling framework proposed by Ngai and Roland,93 in which mobility in the faster component propagates through intermolecular interactions to accelerate relaxation of the slower matrix.
Overall, these findings provide a molecular-level framework for understanding dynamic asymmetry in polymer blends. We note that our results pertain to amorphous systems of syndiotactic PMMA. Given the similar $T_{\text{g}}$ and segmental dynamics of atactic PMMA, we expect our findings to be broadly transferable to amorphous blends containing atactic PMMA as well. The same analytical approaches could be applied to systems with crystallinity or other tacticities to probe how packing and morphology shape asymmetric dynamics. Similar mechanisms are anticipated in other flexible–rigid polymer pairs such as PDMS/PMMA or PDMS/PS, offering a basis to test the generality of this physical picture. More broadly, the results suggest that the influence of one species emerges only at certain local compositions and evolves nonlinearly with environment. This insight may guide strategies for tuning viscoelasticity or transport by adjusting the composition and morphology of dynamically distinct phases.